OpenStudy (anonymous):

write the equation of the line with the given properties in slope intercept form slope=-4, through (-3,6)

OpenStudy (anonymous):

y=mx+b m is your slope and plug in the x and y coordinates and solve for b

OpenStudy (anonymous):

y = mx+b -3 is y and 6 is x solve

OpenStudy (anonymous):

b=-21?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

i actually think its -6?

OpenStudy (anonymous):

how did u get to -6 @lewopz

OpenStudy (anonymous):

yes lew is correct my mind is beat tonight. y = mx+b 6 = -4(-3)+b 6 = 12 + b 6-12 = B -6 = b

OpenStudy (anonymous):

there you go ken i corrected my errors and that is how lew got -6

OpenStudy (anonymous):

Yea ^^ and feels aha

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

now you just plug all numbers back into your equation with out the coordinate to get the equation of your linear function

OpenStudy (anonymous):

no problem.